Copy and complete this table

Income. C. S. MPS. MPC. APC. APS

0. 400. a. - - b c

1000. d. e. f. 0.6. g. h

1500. I. j. k l. m. 0.89


1
Expert's answer
2021-07-07T09:06:48-0400

Income equals consumption plus savings. So at zero income, when consumption is 400, savings equals -400. 

"MPS =\frac{ ∆S}{∆Y}\\

MPC =\frac{ ∆C}{∆Y}\\

MPC +MPS = 1\\

APC =\frac{ C}{I}\\

APS =\frac{ S}{I}\\

APC +APS = 1"

Using these formulas we can complete the table as follows;


"a: 0 - 400= -400\\

b: \frac{C}{I} =\frac{ 400}{0} = infinity\\

c: \frac{S}{I} =\frac{ -400}{0 }= infinity\\

d: MPC × ∆Y = ∆C; 0.6×1000 = 600. Therefore C= 600+400 = 1000\\

e: 1000 - 1000 = 0\\

f: \frac{∆S}{∆Y} =\frac{ 400}{1000} = 0.4\\

g: \frac{C}{I} = \frac{1000}{1000} = 1\\

h:\frac{S}{I} = \frac{0}{1000} = 0\\

j: APS × I = 0.89×1500 = 1335\\

i: 1500-1335 = 165\\

k: \frac{∆S}{∆Y} = \frac{1335}{500} = 2.67\\

l:\frac{ ∆C}{∆Y }= \frac{-835}{500 }= -1.67\\

m: \frac{C}{I} = \frac{165}{1500} = 0.11"